A new printer can run off 1,500 workbooks in 8 hours, while an older printer takes 12 hours to run the same load. If both run at the same time to finish 1,500 workbooks, how many hours would it take.
Answer by Alan3354(69443)

A new printer can run off 1,500 workbooks in 8 hours, while an older printer takes 12 hours to run the same load. If both run at the same time to finish 1,500 workbooks, how many hours would it take.
----------------
Call 1500 the job.
The new printer does 1/8 of the job per hour.
The old one does 1/12 of the job per hour.
Together they do 1/8 + 1/12 per hour = 3/24 + 2/24
= 5/24 job per hour.
--> 24/5 hours to to the job
= 4.8 hours
--------------
There's a shortcut: product over sum
= 12*8/(12+8) = 96/20
= 4.8 hours